OpenStudy (anonymous):

What are the solutions of 6x2 − x − 15 = 0?

OpenStudy (anonymous):

6 x^2 = x+15 (2 x+3) (3 x-5) = 0

OpenStudy (anonymous):

6 (x-1/12)^2-361/24 = 0

OpenStudy (anonymous):

x = -3/2

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

6x(2)-x-15=0, subtract x from 6x, 5x(2)-15=0, add 15 to each side, 5x(2)=15, divide each side by 2, 5x=7.5, divide each side by 5 to get x by itself, x=1.5